OpenStudy (anonymous):

Graph each polynomial f(x)=(x-2)(x+1)(x+3)

OpenStudy (anonymous):

It cuts x-axis at x=-3,-1,2 It cuts y-axis at f(x)=(0-2)(0+1)(0+3)=-6 f(x)=(x-2)(x^2+4x+3)=x^3+2x^2-5x-6 \[f \prime \left( x \right)=3x ^{2}+4x-5\] Tangent is parallel to x-axis where \[3x ^{2}+4x-5=0,x=\frac{ -4\pm \sqrt{16+60} }{ 6 }\] \[x=\frac{ -4\pm2\sqrt{19} }{6}=\frac{ -2\pm \sqrt{19} }{3 }\] f(x) increases where f(x) >0 \[x<\frac{ -2-\sqrt{19} }{ 3 } and x>\frac{ -2+\sqrt{19} }{3}\] f(x) decreases where f(x)<0, or \[\frac{ -2-\sqrt{19} }{3 }<x<\frac{ -2+\sqrt{19} }{3 }\] \[f \prime \prime \left( x \right)=6x+4\] \[at x=\frac{ -2-\sqrt{19} }{3 },f \prime \prime \left( x \right)<0\] therefore there is maxima at this point. \[at x=\frac{ -2+\sqrt{19} }{ 3 },f \prime \prime \left( x \right)>0\] hence there is minima at this point. find maximum and minimum f(x) at these points. and then draw keeping in mind above points.